The heat of a certain type of cooking oil is 1.75cal/(g. Degrees C) how much heat energy is needed to raise the temperature of 2.55 kg of this oil from 23 degrees C to 191 degrees C?
Answer #1

Given that, Specific heat copacity = 1-75 cal / 19.6) change in temperature (AT) = Ty - T, = 191–23=1686 - massag= 2.55 kg =